# EnvInjection: Environmental Prompt Injection Attack to Multi-modal Web Agents

### v2

@@ Abstract
Multi-modal large language model (MLLM)-based web agents interact with webpage environments by generating actions based on screenshots of the webpages. Environmental prompt injection attacks manipulate the environment to induce the web agent to perform a specific, attacker-chosen action–denoted as the target action. However, existing attacks suffer from limited effectiveness or stealthiness, or are impractical in real-world settings. In this work, we propose EnvInjection, a new attack that addresses these limitations. Our attack adds a perturbation to the raw pixel values of the rendered webpage. After these perturbed pixels are mapped into a screenshot, the perturbation induces the web agent to perform the target action. We formulate the task of finding the perturbation as an optimization problem. A key challenge in solving this problem is that the mapping between raw pixel values and screenshot is non-differentiable, making it difficult to backpropagate gradients to the perturbation. To overcome this, we train a neural network to approximate the mapping and apply projected gradient descent to solve the reformulated optimization problem. Extensive evaluation on multiple webpage datasets shows that EnvInjection is highly effective and significantly outperforms existing baselines.
EnvInjection: Environmental Prompt Injection Attack
to Multi-modal Web Agents
@@ 1 Introduction
Figure 1: Illustration of EnvInjection.
@@
However, despite the advanced capabilities of MLLM-based web agents, they remain vulnerable to emerging security and safety threats. One such threat is the environmental prompt injection attack Liao et al. (2025); Zhang et al. (2024); Aichberger et al. (2025); Zhao et al. (2025); Wu et al. (2025), in which an adversary manipulates the environment to induce the agent to perform a specific, attacker-chosen action—referred to as the target action—such as clicking on a designated coordinate on the monitor. This type of attack poses a serious security risk, potentially resulting in consequences such as click fraud, malware downloads, or disclosure of sensitive information.
@@
Environmental prompt injection attacks can be categorized into two types: 1) Webpage-based attacks Liao et al. (2025); Zhang et al. (2024); Xu et al. (2024). These attacks aim to mislead the web agent into generating a target action by modifying a webpage’s source code–for example, by injecting deceptive HTML elements such as pop-up windows. However, most existing webpage-based attacks are heuristic-driven and often exhibit suboptimal effectiveness. Furthermore, they lack stealth or, when stealth is preserved, sacrifice a certain degree of effectiveness, as the injected elements are typically visible to users and easily detected. 2) Screenshot-based attacks Aichberger et al. (2025); Zhao et al. (2025). These attacks add visual perturbations directly to the screenshot of a webpage to increase the likelihood that the web agent performs the target action. However, such attacks are impractical in real-world scenarios, since attackers do not have direct access to modify screenshots, which are captured locally on the user’s device. Furthermore, none of them has discussed the webpage-to-screenshot mapping, reflecting a lack of consideration for this critical aspect. Therefore, while one might attempt to implement these perturbations by modifying the raw pixel values via changes to the webpage’s source code, this approach fails entirely due to the nontrivial webpage-to-screenshot mapping, as demonstrated in our experiments. More details on related work are shown in Section 6.
@@
In this work, we introduce a new webpage-based attack, EnvInjection, which achieves both effectiveness and stealthiness while maintaining practical feasibility. Fig. 1 provides a brief illustration of EnvInjection: the attacker introduces a perturbation to a webpage’s raw pixel values via modifying its source code; this indirectly perturbs the resulting screenshot, thereby misleading the web agent into generating the target action. In particular, to ensure both effectiveness and stealthiness, we formulate the task of finding the perturbation as an optimization problem. The objective is to maximize the probability that the MLLM generates the desired target action (effectiveness), while the constraint bounds the $\ell_{\infty}$-norm of the perturbation to ensure it remains imperceptible to users (stealthiness). Furthermore, since the webpage-to-screenshot mapping varies across monitors, we constrain the perturbation to lie within the overlapping region shared by multiple types of monitors, thereby crafting a universal perturbation.
@@
We conduct an in-depth evaluation of our attack. We begin by constructing extensive datasets of webpages, including synthetic and real webpages. Our extensive evaluation demonstrates that EnvInjection is highly successful and significantly outperforms existing attacks. Specifically, when the web agent uses the MLLM Gemma-3 Team et al. (2025), the success rate of our attack is 0.910 higher than the best-performing baseline. We also perform a series of ablation studies to examine the impact of the number of monitors, perturbation bounds, different categories of prompts, and various target actions. These studies further demonstrate the generalizability of EnvInjection across different configurations and variations.
@@ 3 Threat Model
4 EnvInjection
Our attack EnvInjection aims to achieve both effectiveness and stealthiness by modifying the source code of the target webpage $\omega$. To this end, the attack first introduces a human-imperceptible perturbation $\delta$ to the rendered raw pixel values $I(\omega,d)$ of the target webpage, resulting in modified pixels $I(\omega,d)+\delta$. The attack then implements this perturbation by modifying the source code $\omega$ to obtain a new version $\omega^{\prime}$ such that $I(\omega^{\prime},d)=I(\omega,d)+\delta$. In the following, we first formulate the task of finding the perturbation $\delta$ as an optimization problem, then present our algorithm to solve it, and finally describe how the perturbation is implemented via modifying the source code $\omega$.
@@ 5.1 Experimental Setup
Agent Naive Context Ignoring Fake Completion Combined Screenshot-based EnvInjection
@@ 5.2 Experimental Results
EnvInjection achieves both stealthiness and effectiveness goals and outperforms existing attacks. Table 1 reports the ASR of various attacks averaged across our 10 webpage datasets for different MLLM-based web agents. A detailed breakdown of ASR results for each dataset is provided in Tables 5-9 in Appendix. We observe that EnvInjection consistently achieves high effectiveness and significantly outperforms all baseline attacks. For example, when the web agent uses the MLLM Gemma-3, the highest ASR achieved by existing webpage-based attacks is 0.062, while screenshot-based attacks yield an ASR of 0.000. In contrast, EnvInjection achieves an ASR of 0.972. This substantial improvement stems from the optimization-based nature of EnvInjection, which directly maximizes the likelihood that the agent generates the target action. In comparison, existing webpage-based attacks rely on heuristic injection strategies, and screenshot-based attacks fail to consider the critical webpage-to-screenshot mapping.
@@
Figure 2: Impact of the number of target monitors and $\epsilon$ on the average ASR of EnvInjection across five agents.
@@
Impact of the number of target monitors. Fig. 2 shows the impact of the number of target monitors on the average ASR of our EnvInjection across the five web agents. A detailed breakdown of ASR per dataset is provided in Fig. 11-12 in the Appendix. We observe that ASR decreases slightly as the number of target monitors increases. This is because the perturbation space to be optimized becomes smaller, since we only optimize the perturbation within the overlapping region. Nevertheless, selecting more target monitors enables the attacker to successfully compromise a greater number of users who use different monitors, although the probability of successfully attacking each user decreases slightly on average. Additionally, as shown in Table 1, although webpage-based and screenshot-based attacks are not affected by the number of target monitors, they still perform significantly worse than EnvInjection when the number of target monitors increases.
@@
Impact of the perturbation bound $\epsilon$. Fig. 2 shows the impact of $\epsilon$ on the average ASR of our EnvInjection across the five web agents. A detailed breakdown of ASR per dataset is provided in Fig. 13-14 in the Appendix. We observe that as $\epsilon$ increases from $4/255$ to $32/255$, the ASR rises to nearly 1. This is because a larger $\epsilon$ provides a greater space for optimization. This result further illustrates that our EnvInjection can successfully achieve both effectiveness and stealthiness goals. Note that $\epsilon\leq 16/255$ is generally considered stealthy in prior works Qi et al. (2024); Luo et al. (2024). Examples of the perturbed webpages under different $\epsilon$ are shown in Fig. 5 in Appendix.
@@
User prompts are semantically equivalent variants of the target prompts. Table 11 in Appendix shows the ASR of EnvInjection across different agents when user-specified prompts are semantically equivalent variants of the target prompts but not textually identical. Specifically, ASR is computed by replacing the target prompt $p_{\omega}$ with its semantically equivalent user prompt in Equation 5.1. Given a target prompt, we generate its semantic equivalent user prompt using GPT-4-Turbo OpenAI (2023), guided by the instruction shown in Fig. 8 in Appendix. We observe that even though EnvInjection is not directly optimized for user prompts, it still achieves comparable ASR. For example, for the Gemma-3 agent on the synthetic blog webpage dataset, the ASR using user prompts is 0.957, which is close to the ASR using target prompts, 0.988. This result highlights that EnvInjection can extend to a wide range of user prompts, as long as the user prompt is semantically similar to the target prompt used in optimization.
@@
Other target actions. In our prior experiments, we use click((x,y)) as a target action. Table 10 in Appendix shows the ASR of EnvInjection for other target actions on the synthetic Blog dataset when using Phi-4 Abouelenin et al. (2025) as the MLLM. The results show that our EnvInjection is also highly successful at misleading the web agent to generate other target actions.
@@ 6 Related Work
Environmental prompt injection attacks. Prompt injection attacks have also been extended to web agents. The pop-up attack Zhang et al. (2024) deceives web agents by injecting a misleading pop-up window. EIA Liao et al. (2025) injects HTML elements that are similar to attacker-chosen legitimate elements, thereby tricking the agent into interacting with the injected elements instead of the originals. Screenshot-based attacks Aichberger et al. (2025); Zhao et al. (2025) employ adversarial example techniques Szegedy et al. (2014) to optimize stealthy visual perturbations added to screenshots, thereby maximizing the probability that web agents generate the target action. As discussed in Section 1, unlike prior environmental prompt injection attacks, EnvInjection optimizes perturbations that can be directly implemented by modifying the webpage’s source code, making the attack effective, stealthy, and practical.
@@ 7 Conclusion
In this paper, we propose EnvInjection, the first effective, stealthy, and practical environmental prompt injection attack against MLLM-based web agents. Our attack optimizes a universal perturbation for a target webpage across diverse target monitors, maximizing the probability that web agents perform the target action. Extensive experiments show that our attack largely outperforms baselines.
@@ 8 Limitations
Potential defenses for EnvInjection include analyzing the webpage source code to identify injected or abnormal code snippets, detecting perturbations in screenshots using adversarial example detection methods Carlini and Wagner (2017), and fine-tuning an MLLM through adversarial training Madry et al. (2018) to enhance its robustness against such perturbations. We note that prompt-injection detection methods such as DataSentinel Liu et al. (2025) are not applicable in our setting, as EnvInjection does not rely on injecting explicit textual prompts.

### v3

# WebInject: Prompt Injection Attack to Web Agents

@@ Abstract
Multi-modal large language model (MLLM)-based web agents interact with webpage environments by generating actions based on screenshots of the webpages. In this work, we propose WebInject, a prompt injection attack that manipulates the webpage environment to induce a web agent to perform an attacker-specified action. Our attack adds a perturbation to the raw pixel values of the rendered webpage. After these perturbed pixels are mapped into a screenshot, the perturbation induces the web agent to perform the attacker-specified action. We formulate the task of finding the perturbation as an optimization problem. A key challenge in solving this problem is that the mapping between raw pixel values and screenshot is non-differentiable, making it difficult to backpropagate gradients to the perturbation. To overcome this, we train a neural network to approximate the mapping and apply projected gradient descent to solve the reformulated optimization problem. Extensive evaluation on multiple datasets shows that WebInject is highly effective and significantly outperforms baselines.
WebInject: Prompt Injection Attack to Web Agents
@@ 1 Introduction
Figure 1: Illustration of WebInject.
@@
However, despite the advanced capabilities of MLLM-based web agents, they remain vulnerable to emerging security and safety threats. One such threat is prompt injection attack Liu et al. (2024); Liao et al. (2025); Zhang et al. (2024); Aichberger et al. (2025); Zhao et al. (2025); Wu et al. (2025), in which an adversary manipulates the web environment to induce the agent to perform a specific, attacker-chosen action—referred to as the target action—such as clicking on a designated coordinate on the monitor. This type of attack poses a serious security risk, potentially resulting in consequences such as click fraud, malware downloads, or disclosure of sensitive information.
@@
Prompt injection attacks to web agents can be categorized into two types: 1) Webpage-based attacks Liao et al. (2025); Zhang et al. (2024); Xu et al. (2024). These attacks aim to mislead the web agent into generating a target action by modifying a webpage’s source code–for example, by injecting deceptive HTML elements such as pop-up windows. However, most existing webpage-based attacks are heuristic-driven and often exhibit suboptimal effectiveness. Furthermore, they lack stealth or, when stealth is preserved, sacrifice a certain degree of effectiveness, as the injected elements are typically visible to users and easily detected. 2) Screenshot-based attacks Aichberger et al. (2025); Zhao et al. (2025). These attacks add visual perturbations directly to the screenshot of a webpage to increase the likelihood that the web agent performs the target action. However, such attacks are impractical in real-world scenarios, since attackers do not have direct access to modify screenshots, which are captured locally on the user’s device. Furthermore, none of them has discussed the webpage-to-screenshot mapping, reflecting a lack of consideration for this critical aspect. Therefore, while one might attempt to implement these perturbations by modifying the raw pixel values via changes to the webpage’s source code, this approach fails entirely due to the nontrivial webpage-to-screenshot mapping, as demonstrated in our experiments. More details on related work are shown in Section 6.
@@
In this work, we introduce a new webpage-based attack, WebInject, which achieves both effectiveness and stealthiness while maintaining practical feasibility. Fig. 1 provides a brief illustration of WebInject: the attacker introduces a perturbation to a webpage’s raw pixel values via modifying its source code; this indirectly perturbs the resulting screenshot, thereby misleading the web agent into generating the target action. In particular, to ensure both effectiveness and stealthiness, we formulate the task of finding the perturbation as an optimization problem. The objective is to maximize the probability that the MLLM generates the desired target action (effectiveness), while the constraint bounds the $\ell_{\infty}$-norm of the perturbation to ensure it remains imperceptible to users (stealthiness). Furthermore, since the webpage-to-screenshot mapping varies across monitors, we constrain the perturbation to lie within the overlapping region shared by multiple types of monitors, thereby crafting a universal perturbation.
@@
We conduct an in-depth evaluation of our attack. We begin by constructing extensive datasets of webpages, including synthetic and real webpages. Our extensive evaluation demonstrates that WebInject is highly successful and significantly outperforms existing attacks. Specifically, when the web agent uses the MLLM Gemma-3 Team et al. (2025), the success rate of our attack is 0.910 higher than the best-performing baseline. We also perform ablation studies to examine the impact of the number of monitors, perturbation bounds, different categories of prompts, and various target actions. These studies further demonstrate the generalizability of WebInject across configurations and variations.
@@ 3 Threat Model
4 WebInject
Our attack WebInject aims to achieve both effectiveness and stealthiness by modifying the source code of the target webpage $\omega$. To this end, the attack first introduces a human-imperceptible perturbation $\delta$ to the rendered raw pixel values $I(\omega,d)$ of the target webpage, resulting in modified pixels $I(\omega,d)+\delta$. The attack then implements this perturbation by modifying the source code $\omega$ to obtain a new version $\omega^{\prime}$ such that $I(\omega^{\prime},d)=I(\omega,d)+\delta$. In the following, we first formulate the task of finding the perturbation $\delta$ as an optimization problem, then present our algorithm to solve it, and finally describe how the perturbation is implemented via modifying the source code $\omega$.
@@ 5.1 Experimental Setup
Agent Naive Context Ignoring Fake Completion Combined Screenshot-based WebInject
@@ 5.2 Experimental Results
WebInject achieves both stealthiness and effectiveness goals and outperforms existing attacks. Table 1 reports the ASR of various attacks averaged across our 10 webpage datasets for different MLLM-based web agents. A detailed breakdown of ASR results for each dataset is provided in Tables 5-9 in Appendix. We observe that WebInject consistently achieves high effectiveness and significantly outperforms all baseline attacks. For example, when the web agent uses the MLLM Gemma-3, the highest ASR achieved by existing webpage-based attacks is 0.062, while screenshot-based attacks yield an ASR of 0.000. In contrast, WebInject achieves an ASR of 0.972. This substantial improvement stems from the optimization-based nature of WebInject, which directly maximizes the likelihood that the agent generates the target action. In comparison, existing webpage-based attacks rely on heuristic injection strategies, and screenshot-based attacks fail to consider the critical webpage-to-screenshot mapping.
@@
Figure 2: Impact of the number of target monitors and $\epsilon$ on the average ASR of WebInject across five agents.
@@
Impact of the number of target monitors. Fig. 2 shows the impact of the number of target monitors on the average ASR of our WebInject across the five web agents. A detailed breakdown of ASR per dataset is provided in Fig. 11-12 in the Appendix. We observe that ASR decreases slightly as the number of target monitors increases. This is because the perturbation space to be optimized becomes smaller, since we only optimize the perturbation within the overlapping region. Nevertheless, selecting more target monitors enables the attacker to successfully compromise a greater number of users who use different monitors, although the probability of successfully attacking each user decreases slightly on average. Additionally, as shown in Table 1, although webpage-based and screenshot-based attacks are not affected by the number of target monitors, they still perform significantly worse than WebInject when the number of target monitors increases.
@@
Impact of the perturbation bound $\epsilon$. Fig. 2 shows the impact of $\epsilon$ on the average ASR of our WebInject across the five web agents. A detailed breakdown of ASR per dataset is provided in Fig. 13-14 in the Appendix. We observe that as $\epsilon$ increases from $4/255$ to $32/255$, the ASR rises to nearly 1. This is because a larger $\epsilon$ provides a greater space for optimization. This result further illustrates that our WebInject can successfully achieve both effectiveness and stealthiness goals. Note that $\epsilon\leq 16/255$ is generally considered stealthy in prior works Qi et al. (2024); Luo et al. (2024). Examples of the perturbed webpages under different $\epsilon$ are shown in Fig. 5 in Appendix.
@@
User prompts are semantically equivalent variants of the target prompts. Table 11 in Appendix shows the ASR of WebInject across different agents when user-specified prompts are semantically equivalent variants of the target prompts but not textually identical. Specifically, ASR is computed by replacing the target prompt $p_{\omega}$ with its semantically equivalent user prompt in Equation 5.1. Given a target prompt, we generate its semantic equivalent user prompt using GPT-4-Turbo OpenAI (2023), guided by the instruction shown in Fig. 8 in Appendix. We observe that even though WebInject is not directly optimized for user prompts, it still achieves comparable ASR. For example, for the Gemma-3 agent on the synthetic blog webpage dataset, the ASR using user prompts is 0.957, which is close to the ASR using target prompts, 0.988. This result highlights that WebInject can extend to a wide range of user prompts, as long as the user prompt is semantically similar to the target prompt used in optimization.
@@
Other target actions. In our prior experiments, we use click((x,y)) as a target action. Table 10 in Appendix shows the ASR of WebInject for other target actions on the synthetic Blog dataset when using Phi-4 Abouelenin et al. (2025) as the MLLM. The results show that our WebInject is also highly successful at misleading the web agent to generate other target actions.
@@ 6 Related Work
Prompt injection attacks to web agents. Prompt injection attacks have also been extended to web agents. The pop-up attack Zhang et al. (2024) deceives web agents by injecting a misleading pop-up window. EIA Liao et al. (2025) injects HTML elements that are similar to attacker-chosen legitimate elements, thereby tricking the agent into interacting with the injected elements instead of the originals. Screenshot-based attacks Aichberger et al. (2025); Zhao et al. (2025) employ adversarial example techniques Szegedy et al. (2014) to optimize stealthy visual perturbations added to screenshots, thereby maximizing the probability that web agents generate the target action. As discussed in Section 1, unlike prior prompt injection attacks, WebInject optimizes perturbations that can be directly implemented by modifying the webpage’s source code, making the attack effective, stealthy, and practical.
@@ 7 Conclusion
In this paper, we propose WebInject, the first effective, stealthy, and practical prompt injection attack to web agents. Our WebInject optimizes a universal perturbation for a target webpage across diverse target monitors, maximizing the probability that web agents perform the attacker-chosen target action. Extensive experiments show that our attack largely outperforms baselines.
@@ 8 Limitations
Potential defenses for WebInject include analyzing the webpage source code to identify injected or abnormal code snippets, detecting perturbations in screenshots using adversarial example detection methods Carlini and Wagner (2017), and fine-tuning an MLLM through adversarial training Madry et al. (2018) to enhance its robustness against such perturbations. We note that prompt-injection detection methods such as DataSentinel Liu et al. (2025) are not applicable in our setting, as WebInject does not rely on injecting explicit textual prompts.